OpenStudy (anonymous):

Two objects attract each other with a gravitational force of 50 newtons from a given distance. If the distance between the two objects is increased by 5 times, what will be the changed force of attraction between them?

OpenStudy (anonymous):

\[\text{gravitational force} \propto \dfrac{1}{r^2}\]Do you understand what the above means?

OpenStudy (anonymous):

As F∝m1m1/r^2, F=50N initially as r=1. If r is increased by 5 times, then, r=5 and r^2=25. So, F=50/25N=2N. When the distance between the objects were increased 5 times, the force of attraction got decreased to 2 Newton.